9.Find the equation of a circle with center (1, -1) and that passes through the origin.

10. Find the equation of a circle with center (2, 1) that is tangent to the x-axis.

Answer by stanbon(75887)

Find the equation of a circle with center (1, -1) and that passes through the origin.
Radius = distance from origin to (1,-1) = sqrt(1^2+1^2) = sqrt(2)
--------------------
Ans: x^2 + y^2 = 2
--------------------------

10. Find the equation of a circle with center (2, 1) that is tangent to the x-axis.
The point of tangency is (2,0)
Radius = sqrt[(2-2)^2+(1-0)^2] = 1
------
Ans: (x-2)^2 + (y-1)^2 = 1
------------------------------
Cheers,
Stan H.
--------------